OpenStudy (anonymous):

The function t(x) = 3x + 5 determines how many cans of corn kernels a food truck needs to stock, where x is the number of shifts the crew is going to work in the truck. The crew uses c(t(x)) to find the amount of money to spend on corn. The function c(x) = 2x + 3. Solve for how much money must be spent when the crew is going to work 3 shifts.

OpenStudy (anonymous):

t(x) = 3x + 5 t(x) = 3(3) + 5 t(x) = 9 + 5 t(x) = 14 Is that correct?

OpenStudy (boldjon):

c(x) = 2x+3 x = t(x) = 3x+5 2(3x+5)+3 = 6x+10+3 = 6x+13 x = 3 shifts 6(3)+10+3 = 18+10+3 = 31

OpenStudy (boldjon):

or you can let x = 3 in t(x), and put that number into c(x)

OpenStudy (boldjon):

t(3) = 3(3)+5 = 14 c(14) = 2(14)+3 = 31

OpenStudy (boldjon):

The above may be easier for you

OpenStudy (anonymous):

Then you have to plus that in for c(x) so... c(x) = 2x + 3 c(x) = 2(14) + 3 c(x) = 38 + 3 so c(x) = 31

OpenStudy (mathmale):

The key to understanding what to do here is to understand the notation c(t(x)). Please explain in words what c(t(x)) means.

OpenStudy (anonymous):

Never mind, you already answered lol